OpenStudy (anonymous):

Translate this problem to an equation and solve. The area of a swimming pool is 200 square feet. The length of the pool is 10 feet longer than the width. What is the width of the swimming pool? ?feet

OpenStudy (anonymous):

@JessicaBBY @JaziBoo @Jpenrod @jigglypuff314 @jojo4eva

OpenStudy (anonymous):

the area of a swimming pool is 200 square feet. the lengthof the pool is 10 feet longer than the width.what is the width of the swimming pool? . Let w = width of pool then w+10 = length of pool . w(w+10) = 200 w^2 + 10w = 200 w^2 + 10w - 200 = 0 (w+20)(w-10) = 0 w = {-20, 10} you can throw out the negative solution (extraneous) leaving: w = 10 feet

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (jojo4eva):

ya because the width of the pool would be 20 since 20 times 10=200

OpenStudy (anonymous):

Translate this problem to an equation and solve. The area of a patio is 72 square feet. The length of the patio is 1 foot longer than the width. What is the width of the patio? feet

OpenStudy (phi):

This patio problem is the same kind as the swimming pool problem Can you give it a try ? First, do you know a formula for the area of a rectangle ? can you write it down ?

OpenStudy (anonymous):

Area = ½ × b × h b = base h = vertical height

OpenStudy (phi):

that is the area of a triangle (a triangle is ½ of a rectangle) so make a note: review formulas for area of simple figures. A= b*h b= base (or length) h= height (or width)

OpenStudy (anonymous):

@ryanvarghese12790

OpenStudy (phi):

now we look at the info they give us The area of a patio is 72 square feet. in A = L*W (let's use length times width) where does the 72 go ?

OpenStudy (anonymous):

AREA

OpenStudy (phi):

yes, but write down the equation, with the 72 replacing the A 72 = L * W

OpenStudy (anonymous):

1 ft is the length im guessing

OpenStudy (phi):

The length of the patio is 1 foot longer than the width. You have to interpret this statement. we do it this way.. toss out irrelevant words... (like "the", etc) length is 1 foot longer than width change length to L (we use letters in algebra) L is "is" almost always means = L = 1 foot longer if you see 1 more than, or 1 bigger than, or 1 longer than, they get translated into 1+ L = 1+ width change width to W L = 1+W

OpenStudy (phi):

so far we have 72 = L * W L = 1+W What is the width of the swimming pool? That means you want to find W. start with 72=L * W replace the L with (1+W) because the 2nd equation tells you L= (1+W) (notice we need the parens) 72= (1+W)W can you solve this? or do you need more help?

OpenStudy (anonymous):

i sort of get it but then i dont get it

OpenStudy (phi):

The ideas do make sense, so you have to figure out what is confusing you, and get rid of the "bad idea" that keeps you from understanding.

OpenStudy (phi):

72= (1+W)W can you solve this? or do you need more help?

OpenStudy (anonymous):

i need help but im get off right now but thanks for the help

OpenStudy (phi):

The equation 72 = (1+W)W is a quadratic... if you "distribute" the W you get 72 = W + W^2 rearrange into standard form you get W^2 + W - 72 =0 you can factor this to find the two values of W. Only 1 is positive so that will be the answer. See http://www.khanacademy.org/math/algebra/quadratics/quadratic_odds_ends/v/solving-a-quadratic-by-factoring for how to factor.